1) Which circuit has the largest maximum current?


2) In which circuit is the most power supplied by the AC generator?


- A conceptual and numerical explaination would be appreciated!


Circuit A) A 10 Ohm resistor B) 0.1 microFarad capacitor C) 1mH inductor









1) Which circuit has the largest maximum current?
Answer #1

1.

1st circuit..... i = v/r = 120/ 10 = 12 A

2nd circuit.....i= v/Xc = V / 1/(2?fC) = 120 x 2x 3.14 x 60x 0.1 x 10^-6 = 4.52 x 10^-3 A

3rd circuit....i = V x XL = 120 x 2x 3.14 x 60x 10^-3 = 45.23 A


so, 3rd circuit has the largest maximum current


and as power P is directly proportional to i^2.... 3rd circuit is the most power supplied by the AC generator
Answer #2

in 1st circuit R= 10 ohm


in 2nd circuit X = 1/(2?fC) = 26525.8 ohm


in 3rd circuit X = 2*3.14 x 60x 10^-3 = 2.65 ohm



resistance in 3rd circuit is minimum so current in 3rd circuit is maximum.

3rd circuit is the most power supplied by the AC generator
Answer #3

Power in 1st circuit=V^2/R=120^2/10=1440W

Power in 2nd circuit=V^2/Xc=V^2*6.28*f*c=0.54W

Power in 3rd circuit=V^2/Xl=V^2/6.28*f*L=382165.60W

So largest power in 3rd circuit.(answer)

Now for maximum current,circuit with lowest resistance

R1=10,R2=26539,R3=0.3768

So maximum current in 3rd circuit.(answer)